\begin{thebibliography}{99}

\bibitem{AQVV} G.D.\,Anderson, S.L.\,Qiu, M.K.\,Vamanamurthy and M.\,Vuorinen,
Generalized elliptic integrals and modular equations,
\textit{Pacific J.\,Math.} \textbf{192} (2000), no. 1, 1--37. 

\bibitem{AAR}
G.\,Andrews, R.\,Askey and R.\,Roy,
\textit{Special functions},
Encyclopedia of Mathematics and its Applications, 71. 
Cambridge University Press, Cambridge, 1999.

% \bibitem{B} N.\,K.\,Bari,
% Biorthogonal systems and bases in Hilbert space. (Russian) 
% Moskov.\,Gos.\,Univ.\,U\u{e}nye Zapiski Matematika \textbf{148} (4), 
% (1951), 69-107.
% \bibitem{BF} M.S.\,Berger and L.E.\,Fraenkel,
% On the asymptotic solution of a nonlinear Dirichlet problem, 
% \textit{J.\ Math.\ Mech.} \textbf{19} (1969/1970), 553--585. 

%\bibitem{Be} B.C.\,Berndt,
%\textit{Ramanujan's notebooks. Part V},
%Springer-Verlag, New York, 1998.


%\bibitem{BBG} B.C.\,Berndt, S.\,Bhargava and F.G.\,Garvan,
%Ramanujan's theories of elliptic functions to alternative bases,
%\textit{Trans.\,Amer.\,Math.\,Soc.} \textbf{347} (1995), no.\,11, 4163--4244.
 
% \bibitem{BBH} F.\,Bethuel, H.\,Brezis and F.\,Helein,
% 	\textit{Ginzburg-Landau vortices},
% 	Progress in Nonlinear Differential Equations and their 
% 	Applications, 13. Birkhauser Boston, Inc., Boston, MA, 
% 	1994. 

%\bibitem{BL}
%R.\,Bhatia and R.-C.\,Li,
%An interpolating family of means,
%\textit{Commun.\,Stoch.\,Anal.} \textbf{6} (2012), no. 1, 15--31. 

\bibitem{BY}
B.A.\,Bhayo and L.\,Yin,
On generalized $(p,q)$-elliptic integrals,
preprint, arXiv:1507.00031.

%\bibitem{BBCDG} P.\,Binding, L.\,Boulton, J.\,\v{C}epi\v{c}ka, P.\,Dr\'{a}bek and P.\,Girg,
%Basis properties of eigenfunctions of the p-Laplacian, 
%\textit{Proc.\,Amer.\,Math.\,Soc.}\,\textbf{134} (2006), no.\,12, 3487--3494.

% \bibitem{BB} D.\,Borwein and P.\,B.\,Borwein,
% A generalized arithmetic-geometric mean, Problem 83-12 
% in PROBLEMS AND SOLUTIONS, \textit{SIAM Review} \textbf{25} (1983), 401.

%\bibitem{BB} J.M.\,Borwein and P.B.\,Borwein,
%On the mean iteration $(a,b) \leftarrow ((a+3b)/4, (\sqrt{ab}+b)/2)$. 
%\textit{Math.\,Comp.} \textbf{53} (1989), no.\,187, 311--326. 

%\bibitem{BB2} J.M.\,Borwein and P.B.\,Borwein,
%A cubic counterpart of Jacobi's identity and the AGM,
%\textit{Trans.\,Amer.\,Math.\,Soc.} \textbf{323} (1991), no.\,2, 691--701. 

\bibitem{BB3} J.M.\,Borwein and P.B.\,Borwein,
\textit{Pi and the AGM}, 
A study in analytic number theory and computational complexity. 
Reprint of the 1987 original. Canadian Mathematical Society Series 
of Monographs and Advanced Texts, 4. A Wiley-Interscience Publication. 
John Wiley \& Sons, Inc., New York, 1998

%\bibitem{BBG1} J.\,Borwein, P.\,Borwein and F.\,Garvan,
%Hypergeometric analogues of the arithmetic-geometric mean iteration,
%\textit{Constr.\,Approx.} \textbf{9} (1993), no.\,4, 509--523. 
%
%\bibitem{Br} R.P.\,Brent,
%Fast multiple-precision evaluation of elementary functions,
%\textit{J.\,Assoc.\,Comput.\,Mach.} \textbf{23} (1976), no.\,2, 242--251.

%\bibitem{Bu} F.D.\,Burgoyne,
%Generalized trigonometric functions,
%\textit{Math.\,Comp.} \textbf{18} (1964), 314--316. 
%
%\bibitem{BE} P.J.\,Bushell and D.E.\,Edmunds,
%Remarks on generalized trigonometric functions,
%\textit{Rocky Mountain J.\,Math.}\,\textbf{42} (2012), no.\,1, 25--57.
%
%\bibitem{BL} L.\,Boulton and G.\,Lord,
%Approximation properties of the $q$-sine bases, 
%\textit{Proc.\,R.\,Soc.\,Lond.\,Ser.\,A Math.\,Phys.\,Eng.\,Sci.}
%\textbf{467} (2011), no.\,2133, 2690--2711.  

% \bibitem{CI} N.\,Chafee and E.F.\,Infante,
% A bifurcation problem for a nonlinear partial 
% differential equation of parabolic type, 
% \textit{Applicable Anal.} \textbf{4} (1974/75), 17--37.

%\bibitem{Ch} H.H.\,Chan,
%On Ramanujan's cubic transformation formula for ${}_2F_1(\frac13,\frac23;1;z)$,
%\textit{Math.\,Proc.\,Cambridge Philos.\,Soc.} \textbf{124} (1998), no.\,2, 193--204. 

%\bibitem{C} B.\,D.\,Craven, 
%Stone's theorem and completeness of orthogonal systems,
%\textit{J.\,Austral.\,Math.\,Sci.} \textbf{12} (1971), 
%211-223. 

%\bibitem{DEM} M.\,del Pino, M.\,Elgueta, R.\,Man\'{a}sevich,
%A homotopic deformation along $p$ of a Leray-Schauder degree 
%result and existence for $(\vert u'\vert 			
%^{p-2}u')'+f(t,u)=0,\;u(0)=u(T)=0,\;p>1$, 
%\textit{J. Differential Equations} \textbf{80} (1989), 1--13. 

%\bibitem{Do} O.\,Do\v{s}l\'{y},
%Half-linear differential equations,
%\textit{Handbook of differential equations}, 
%161--357, Elsevier/North-Holland, Amsterdam, 2004.

%\bibitem{DoR} O.\,Do\v{s}l\'{y} and P.\,\v{R}eh\'{a}k,
%\textit{Half-linear differential equations}, 
%North-Holland Mathematics Studies, 202. Elsevier Science B.V.,
%Amsterdam, 2005. 

%\bibitem{DKT} P.\,Dr\'{a}bek, P.\,Krej\v{c}\'{i} and 
%P.\,Tak\'{a}\v{c},
%\textit{Nonlinear differential equations}, 
%Papers from the Seminar on Differential Equations held in 
%Chvalatice, June 29--July 3, 1998. Chapman \& Hall/CRC 
%Research Notes in Mathematics, 404. Chapman \& Hall/CRC, 
%Boca Raton, FL, 1999.

\bibitem{DM} P.\,Dr\'{a}bek and R.\,Man\'{a}sevich,
On the closed solution to some nonhomogeneous eigenvalue problems 
with $p$-Laplacian,
\textit{Differential Integral Equations}
\textbf{12} (1999), 773--788. 

\bibitem{Du} P.\,Duren,
The Legendre relation for elliptic integrals,
\textit{Paul Halmos}, 305--315, Springer, New York, 1991.

%\bibitem{EGL} D.E.\,Edmunds, P.\,Gurka and J.\,Lang,
%Properties of generalized trigonometric functions, 
%\textit{J.\,Approx.\,Theory} \textbf{164} (2012), no.\,1, 47--56. 
%
%\bibitem{EGL2} D.E.\,Edmunds, P.\,Gurka and J.\,Lang,
%Basis properties of generalized trigonometric functions,
%\textit{J.\,Math.\,Anal.\,Appl.} \textbf{420} (2014), no.\,2, 1680--1692. 

%\bibitem{E} A.\,Elbert,
%A half-linear second order differential equation. 
%\textit{Qualitative 
%theory of differential equations, Vol. I, II} (Szeged, 1979),
%pp. 153--180, Colloq. Math. Soc. Janos Bolyai, 30,
%North-Holland, Amsterdam-New York, 1981. 

\bibitem{El} E.B.\,Elliott,
A formula including Legendre's $EK'+KE'-KK'=\frac12 \pi$,
\textit{Messenger Math.} \textbf{33} (1903/1904), 31--32.

\bibitem{EMOT} A.\,Erd\'{e}lyi, W.\,Magnus, F.\,Oberhettinger and F.G.\,Tricomi,
\textit{Higher transcendental functions. Vol. I},
Based on notes left by Harry Bateman. With a preface by Mina Rees. 
With a foreword by E. C. Watson. Reprint of the 1953 original. 
Robert E. Krieger Publishing Co., Inc., Melbourne, Fla., 1981.

%\bibitem{EMOT2} A.\,Erd\'{e}lyi, W.\,Magnus, F.\,Oberhettinger and F.G.\,Tricomi,
%\textit{Higher transcendental functions. Vol. II}, 
%Based on notes left by Harry Bateman. Reprint of the 1953 original. Robert E. Krieger Publishing Co., Inc., Melbourne, Fla., 1981.

%\bibitem{G}
%F.\,Garvan,
%Cubic modular identities of Ramanujan, hypergeometric functions and analogues of the arithmetic-geometric mean iteration,
%The Rademacher legacy to mathematics (University Park, PA, 1992), 245--264, 
%\textit{Contemp.\,Math.} \textbf{166}, Amer. Math. Soc., Providence, RI, 1994. 

%\bibitem{GK} I.\,C.\,Gohberg and M.\,G.\,Kre\u{\i}n,
%\textit{Introduction to the theory of linear nonselfadjoint operators}, 
%Translated from the Russian by A. Feinstein. Translations of Mathematical Monographs, Vol. 18 American Mathematical Society, Providence, R.I. 1969

% \bibitem{GV} M.\,Guedda and L.\,V\'eron,
% Bifurcation phenomena associated to the $p$-Laplace operator,
% \textit{Trans.\ Amer.\ Math.\ Soc.} \textbf{310} (1988),
% 419--431.

%\bibitem{H} J.\,R.\,Higgins,
%\textit{Completeness and basis properties of sets of special functions}, 
%Cambridge Tracts in Mathematics, Vol. 72. Cambridge University Press, 
%Cambridge-New York-Melbourne, 1977. 

% \bibitem{J} C.G.J.\,Jacobi, \textit{Fundamenta nova theoriae functionum 
% ellipticarum}, Konigsberg (1829) (in Latin),
% ISBN 978-1-108-05200-9, Reprinted by Cambridge University Press 2012:\\
% \texttt{http://archive.org/details/fundamentanovat00jacogoog}

\bibitem{KT} T.\,Kamiya and S.\,Takeuchi,
Complete $(p,q)$-elliptic integrals with application to a family of means, 
preprint, 
arXiv:1507.01383.

% \bibitem{KLO} P.\,Korman, Y.\,Li and T.\,Ouyang,
% Exact multiplicity results for boundary value problems 
% with nonlinearities generalising cubic, 
% \textit{Proc.\ Roy.\ Soc.\ Edinburgh Sect.\ A}
% \textbf{126} (1996), 599--616.

\bibitem{LE} J.\,Lang and D.E.\,Edmunds,
\textit{Eigenvalues, embeddings and generalised trigonometric functions}, 
Lecture Notes in Mathematics, 2016. Springer, Heidelberg, 2011. 

%\bibitem{Li} P.\,Lindqvist,
%Some remarkable sine and cosine functions,
%\textit{Ricerche Mat.}\,\textbf{44} (1995), no.\,2, 269--290.

%\bibitem{LP} P.\,Lindqvist and J.\,Peetre,
%$p$-arclength of the $q$-circle,
%\textit{The Mathematics Student} \textbf{72} (2003), 139--145.
%
%\bibitem{LP2} P.\,Lindqvist and J.\,Peetre,
%\textit{Comments on Erik Lundberg's 1879 thesis. Especially on the work of 
%G\"{o}ran Dillner and his influence on Lundberg},
%Memorie dell'Instituto Lombardo
%(Classe di Scienze Matem.\,Nat.) 31, 2004.

%\bibitem{M} D.S.\,Mitrinovi\'{c},
%\textit{Analytic inequalities}, 
%In cooperation with P. M. Vasi\'{c},
%Die Grundlehren der mathematischen Wissenschaften, Band 165 Springer-Verlag, New York-Berlin, 1970.

%\bibitem{N} Y.\,Naito,
%Uniqueness of positive solutions of quasilinear differential 
%equations,
%\textit{Differential Integral Equations} \textbf{8} (1995), no. 7, 1813-1822. 

%\bibitem{Na} N.\,Nakamura,
%Order relations among some interpolating families of means,
%\textit{Toyama Math.\,J.} \textbf{35} (2012), 35--48.
 
% \bibitem{O} M.\,\^Otani,
% 	On certain second order ordinary differential equations
% 	associated with Sobolev-Poincar\'{e}-type inequalities,
% 	\textit{Nonlinear Anal.} \textbf{8} (1984), 1255--1270.

%\bibitem{QH} F.\,Qi and Z.\,Huang,
%Inequalities of the complete elliptic integrals, 
%Tamkang J.\,Math.\,\textbf{29} (1998), no.\,3, 165--169. 

%\bibitem{Sa} E.\,Salamin,
%Computation of $\pi$ using arithmetic-geometric mean, 
%\textit{Math.\,Comp.} \textbf{30} (1976), no.\,135, 565--570. 

%\bibitem{Sh} D.\,Shelupsky,
%A generalization of the trigonometric functions, 
%\textit{Amer.\,Math.\,Monthly} \textbf{66} (1959), 879--884. 


%\bibitem{S} I.\,Singer,
%\textit{Bases in Banach spaces. I}, 
%Die Grundlehren der mathematischen Wissenschaften, Band 154. 
%Springer-Verlag, New York-Berlin, 1970. 

\bibitem{T} S.\,Takeuchi,
Generalized Jacobian elliptic functions and their application to bifurcation problems 
associated with $p$-Laplacian, 
\textit{J.\,Math.\,Anal.\,Appl.}\,\textbf{385} (2012), no.\,1, 24--35. 

\bibitem{T2} S.\,Takeuchi,
The basis property of generalized Jacobian elliptic functions,
\textit{Commun.\,Pure Appl.\,Anal.} \textbf{13} (2014), no.\,6, 2675--2692.

\bibitem{T3} S.\,Takeuchi,
A new form of the generalized complete elliptic integrals,
\textit{Kodai Math.\ J.} \textbf{39} (2016), no.\,1, 202--226.

\bibitem{T4} S.\,Takeuchi,
Complete $p$-elliptic integrals and a computation formula of $\pi_p$ for $p=4$,
preprint, arXiv:1503.02394.

% \bibitem{TY} S.\,Takeuchi and Y.\,Yamada,
% Asymptotic properties of a reaction-diffusion equation with 
% degenerate $p$-Laplacian,
% \textit{Nonlinear Anal.} \textbf{42} (2000), 41--61.
% \bibitem{T2} S.\,Takeuchi, in preparation. 
% \bibitem{WK} S-H.\,Wang and N.D.\,Kazarinoff,
% Bifurcation and stability of positive solutions of a 
% two-point boundary value problem,
% \textit{J.\ Austral.\ Math.\ Soc.\ Ser.\ A} 
% \textbf{52} (1992), 334--342. 

%\bibitem{WCJ} M.-K.\,Wang, Y.-M.\,Chu and Y.-P.\,Jiang,
%Ramanujan's cubic transformation and generalized modular equation,
%arXiv preprint, arXiv:1305.6525.
%
%\bibitem{Wa} K.\,Watanabe,
%Planar $p$-elastic curves and related generalized complete elliptic integrals,
%\textit{Kodai Math.\,J.} \textbf{37} (2014),  no.\,2, 453--474.
%
%\bibitem{WW} E.T.\,Whittaker and G.N.\,Watson,
%\textit{A course of modern analysis}, 
%An introduction to the general theory of infinite processes 
%and of analytic functions; with an account of the principal
%transcendental functions. Reprint of the fourth (1927) edition. 
%Cambridge Mathematical Library. Cambridge University Press, 
%Cambridge, 1996.
%

\bibitem{YH} L.\,Yin and L.-G.\,Huang,
Inequalities for the generalized trigonometric and hyperbolic functions 
with two parameters,
\textit{J.\ Nonlinear Sci.\ Appl.} \textbf{8} (2015), no. 4, 315--323.
 

%%\bibitem{Y} K.\,Yosida,
%%\textit{Functional analysis},
%%Reprint of the sixth (1980) edition. Classics in Mathematics. Springer-Verlag, Berlin, 1995.
\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsthm, amsmath, amssymb}

\numberwithin{equation}{section}

\theoremstyle{plain}	     
\newtheorem{thm}{Theorem}[section] 
\newtheorem{cor}[thm]{Corollary}

\newtheorem{prop}[thm]{Proposition}
     
\theoremstyle{definition}

\theoremstyle{remark} 
\newtheorem{rem}[thm]{Remark}
	
\newcommand{\disp}{\displaystyle}

\begin{document}
\title{Legendre-type relations for generalized complete elliptic integrals
\footnote{This work was supported by JSPS KAKENHI Grant Number 24540218.}}
\author{Shingo Takeuchi\\
Department of Mathematical Sciences\\
Shibaura Institute of Technology
\thanks{307 Fukasaku, Minuma-ku,
Saitama-shi, Saitama 337-8570, Japan. \endgraf
{\it E-mail address\/}: shingo@shibaura-it.ac.jp \endgraf
{\it 2010 Mathematics Subject Classification.} 
33E05, 33C75, 34L10}}
\date{}

\maketitle

\begin{abstract}
Legendre's relation for the complete elliptic integrals of the 
first and second kinds is generalized. The proof depends on an application 
of the generalized trigonometric functions and is alternative to
the proof for Elliott's identity.
\end{abstract}

\textbf{Keywords:} 
Legendre's relation,
complete elliptic integrals,
generalized trigonometric functions,
Elliott's identity

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Introduction}

Let $k \in [0,1)$. The complete elliptic integrals of the first kind
$$K(k)=\int_0^1 \frac{dt}{\sqrt{(1-t^2)(1-k^2t^2)}}$$
and of the second kind
$$E(k)=\int_0^1 \sqrt{\frac{1-k^2t^2}{1-t^2}}\,dt$$
play important roles in classical analysis. 
In this paper, we consider generalizations of $K(k)$ and $E(k)$ as
$$K_{p,q,r}(k):=\int_0^1 \frac{dt}{(1-t^q)^{1/p} (1-k^qt^q)^{1/r}}$$
and
$$E_{p,q,r}(k):=\int_0^1 \frac{(1-k^qt^q)^{1/r^*}}{(1-t^q)^{1/p}}\,dt,$$
where $p \in \mathbb{P}^*:=(-\infty,0) \cup (1,\infty],\ 
q,\ r \in (1,\infty)$ and $1/s+1/s^*=1$. 
For $p=\infty$ we regard $K_{p,q,r}$ and $E_{p,q,r}$ as 
$$K_{\infty,q,r}(k):=\int_0^1\frac{dt}{(1-k^qt^q)^{1/r}},\quad
E_{\infty,q,r}(k):=\int_0^1(1-k^qt^q)^{1/r^*}\,dt.$$
Under the convention that $1/\infty=0$ and $1/0=\infty$,
we should note that
$s \in \mathbb{P}^*$ if and only if $s^* \in (0,\infty)$,
particularly, $\infty^*=1$.
In case $p=q=r=2$, $K_{p,q,r}(k)$ and $E_{p,q,r}(k)$ are reduced to 
the classical $K(k)$ and $E(k)$, respectively.

There is a lot of literature about 
the generalized complete elliptic integrals.
$K_{p,q,p}$ is introduced
in \cite{T} with a generalization of the Jacobian elliptic function
with a period of $4K_{p,q,p}$ to study a bifurcation problem of a bistable 
reaction-diffusion equation involving $p$-Laplacian.
%In that work, the constant $\pi_{p/2,p}$ appears for $p>2$, 
%which is equal to $\pi_{s,q}$ with $s=p/2$ defined above.
Relationship between $K_{p,q,p}$ and $E_{p,q,p^*}$ has been observed
in \cite{BY,YH}.
Regarding $K_{p,q,p^*}$,
another generalization of Jacobian elliptic function with a period of $K_{p,q,p^*}$
is given and the basis properties for the family of these functions
are shown in \cite{T2}. Moreover, $K_{p,q,p^*}$ is also applied to a problem 
on Bhatia-Li's mean and a curious relation between $K_{p,q,p^*}$ and $E_{p,q,p^*}$
is given in \cite{KT}. 
%Concerning Legendre's relation, $K_{p,p,p^*}$ and $E_{p,p,p^*}$ satisfy
%have been applied
%in \cite{T3,T4} to generalize \eqref{eq:L} and 
%to establish computation formulas of $\pi_p$
%(only for $p=3$ and $4$), the generalized $\pi$ defined below.

It is well known that $K(k)$ and $E(k)$ satisfy the famous \textit{Legendre's relation}
(see, for example, \cite{AAR, BB3, Du}):
\begin{equation}
\label{eq:L}
E(k)K(k')+K(k)E(k')-K(k)K(k')=\frac{\pi}{2},
\end{equation}
where $k'=\sqrt{1-k^2}$.
Our purpose in the present paper 
is to generalize Legendre's relation \eqref{eq:L}
to the generalized complete elliptic integrals above.

To state the results, we will give some notations. 
For $p \in \mathbb{P}^*$ and $q \in (1,\infty)$, let 
$$\pi_{p,q}:=2\int_0^1 \frac{dt}{(1-t^q)^{1/p}}
=\frac{2}{q}B\left(\frac{1}{q},\frac{1}{p^*}\right),$$
where $B$ denotes the beta function.
In particular, $\pi_{\infty,q}=2$ for any $q \in (1,\infty)$.
We write $K_{p,q}:=K_{p,q,q^*},\ E_{p,q}:=E_{p,q,q^*}$ for $p \in \mathbb{P}^*$
and $q \in (1,\infty)$; $K_p:=K_{p,p,p^*},\ E_p:=E_{p,p,p^*},\ 
\pi_p:=\pi_{p,p}$ for $p \in (1,\infty)$.

\begin{thm}
\label{thm:pqrL}
Let $p \in \mathbb{P}^*,\ q,\ r \in (1,\infty)$ and $k \in (0,1)$. Then
\begin{multline}
\label{eq:pqrL}
E_{p,q,r^*}(k)K_{p,r,q^*}(k')
+K_{p,q,r^*}(k)E_{p,r,q^*}(k')\\
-K_{p,q,r^*}(k)K_{p,r,q^*}(k')
=\frac{\pi_{p,q}\pi_{s,r}}{4},
\end{multline}
where $k':=(1-k^q)^{1/r}$ and $1/s=1/p-1/q$.
\end{thm}

\begin{cor}[Case $q=r$]
\label{cor:q=r}
Let $p \in \mathbb{P}^*,\ q \in (1,\infty)$ and $k \in (0,1)$. Then
\begin{equation}
\label{eq:pqL}
E_{p,q}(k)K_{p,q}(k')
+K_{p,q}(k)E_{p,q}(k')\\
-K_{p,q}(k)K_{p,q}(k')
=\frac{\pi_{p,q}\pi_{s,q}}{4},
\end{equation}
where $k':=(1-k^q)^{1/q}$ and $1/s=1/p-1/q$.
\end{cor}

\begin{cor}[\cite{T3}, Case $p=q=r$]
Let $p \in (1,\infty)$ and $k \in (0,1)$. Then
\begin{equation}
\label{eq:pL}
E_{p}(k)K_{p}(k')
+K_{p}(k)E_{p}(k')\\
-K_{p}(k)K_{p}(k')
=\frac{\pi_{p}}{2},
\end{equation}
where $k':=(1-k^p)^{1/p}$.
\end{cor}

\begin{rem}
Using \eqref{eq:pL}, the author establishes computation formulas of $\pi_p$
for $p=3$ in \cite{T3}; for $p=4$ in \cite{T4}.
\end{rem}

In fact, \eqref{eq:pqrL} is equivalent to \textit{Elliott's identity}
\eqref{eq:elliott} below.
The advantage of our result lies in the facts that  
it is understandable without acknowledge of hypergeometric functions
and that its proof gives an alternative proof for Elliott's identity with straightforward calculations. 
%Our proof of Theorem \ref{thm:pqrL} is, however, 
%not completely satisfactory,
%because it depends upon prior knowledge of the identity.   

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Proof of Theorem \ref{thm:pqrL}}

The following property immediately follows from the definitions of $K_{p,q,r}$ and $E_{p,q,r}$.

\begin{prop}
\label{prop:behaviour}
Let $p \in \mathbb{P}^*,\ q,\ r \in (1,\infty)$.
Then, 
$K_{p,q,r}(k)$ is increasing on $[0,1)$ and
\begin{align*}
K_{p,q,r}(0)&=\frac{\pi_{p,q}}{2},\\
\lim_{k \to 1-0}K_{p,q,r}(k)&=
\begin{cases}
\infty & \mbox{if}\ 1/p+1/r \ge 1,\\
\pi_{u,q}/2\ (1/u=1/p+1/r) & \mbox{if}\ 1/p+1/r<1;
\end{cases}
\end{align*}
and $E_{p,q,r}(k)$ is decreasing on $[0,1]$ and
$$E_{p,q,r}(0)=\frac{\pi_{p,q}}{2},\quad
E_{p,q,r}(1)
=\frac{\pi_{v,q}}{2}\ (1/v=1/p-1/r^*).$$
%\frac{1}{q}B\left(\frac{1}{q},\frac{1}{p^*}+\frac{1}{r^*}\right).$$
\end{prop}

For $p \in \mathbb{P}^*$ and $q \in (1,\infty)$,
the \textit{generalized trigonometric function} $\sin_{p,q}{x}$ is the inverse function of
$$\sin_{p,q}^{-1}{x}:=
\begin{cases}
\disp \int_0^x \frac{dt}{(1-t^q)^{1/p}} & \mbox{if}\ p \neq \infty,\\
x & \mbox{if}\ p=\infty.
\end{cases}
$$
Clearly, $\sin_{p,q}{x}$ is increasing function from $[0,\pi_{p,q}/2]$ onto $[0,1]$.
%,
%where
%$$\pi_{p,q}:=2\sin_{p,q}^{-1}{1}
%=2 \displaystyle \int_0^1 \dfrac{dt}{(1-t^q)^{1/p}}
%=\frac2q B\left(\frac{1}{q},\frac{1}{p^*}\right).$$
%We also define $\pi_{p,q}$ for negative $p$ by the right-hand side of 
%the equation above. Moreover, $\pi_{\pm \infty,q}:=2$.

For $p=q=2$, $\sin_{p,q}{\theta}$ and $\pi_{p,q}=2\sin_{p,q}^{-1}{1}$ are 
identical to the classical $\sin{\theta}$ and $\pi$, respectively.
Moreover, $\sin_{p,q}{\theta}$ and $\pi_{p,q}$ 
play important roles to express the solutions $(\lambda,u)$ 
of inhomogeneous eigenvalue problem of $p$-Laplacian
$-(|u'|^{p-2}u')'=\lambda |u|^{q-2}u,\ p,\ q \in (1,\infty)$,
with a boundary condition (see \cite{DM,LE,T}
and the references given there).

For $p \neq \infty$ and $x \in (0,\pi_{p,q}/2)$, we also define 
$\cos_{p,q}{x}:=(\sin_{p,q}{x})'$.
It is easy to check that for $x \in (0,\pi_{p,q}/2)$,
\begin{gather*}
\cos_{p,q}^p{x}+\sin_{p,q}^q{x}=1,\quad 
%(\sin_{p,q}{x})'=\cos_{p,q}{x},\notag \\
(\cos_{p,q}{x})'=-\frac{q}{p}\sin_{p,q}^{q-1}{x}\cos_{p,q}^{2-p}{x}.
%(\cos_{p,q}^{p-1}{x})'=-\frac{q}{p^*}\sin_{p,q}^{q-1}{x}.
\end{gather*}

Now, we apply the generalized trigonometric function to the generalized
complete elliptic integrals.
For $p \in \mathbb{P}^*$ and $q,\ r \in (1,\infty)$, 
using $\sin_{p,q}{\theta}$ and $\pi_{p,q}$,  
we can express $K_{p,q,r}(k)$ and $E_{p,q,r}(k)$ 
as follows.
\begin{align*}
K_{p,q,r}(k)&=\int_0^{\pi_{p,q}/2}
\frac{d\theta}{(1-k^q\sin_{p,q}^{q}{\theta})^{1/r}},\\
E_{p,q,r}(k)&=\int_0^{\pi_{p,q}/2}
(1-k^q\sin_{p,q}^{q}{\theta})^{1/r^*}\,d\theta.
\end{align*}

Then, we see that 
the functions $K_{p,q,r}(k)$ and $E_{p,q,r}(k)$ satisfy a system of linear 
differential equations.
\begin{prop}
\label{prop:p-differential}
Let $p \in \mathbb{P}^*,\ q,\ r \in (1,\infty)$. Then,
\begin{align*}
\frac{dE_{p,q,r}}{dk}&=\frac{q(E_{p,q,r}-K_{p,q,r})}{r^*k},\\
\frac{dK_{p,q,r}}{dk}&=\dfrac{aE_{p,q,r}-(a-k^q)K_{p,q,r}}{k(1-k^q)},
\end{align*}
where $a:=1+q/r^*-q/p$.
\end{prop}

\begin{proof}
We consider the case $p \neq \infty$.
Differentiating $E_{p,q,r}(k)$ we have
\begin{align*}
\frac{dE_{p,q,r}}{dk}
%&=\int_0^{\pi_{p,q}/2}
%\frac{d}{dk}(1-k^q\sin_{p,q}^q{\theta})^{1/r^*}\,d\theta\\
&=\frac{q}{r^*}\int_0^{\pi_{p,q}/2}
\dfrac{-k^{q-1}\sin_{p,q}^q{\theta}}{(1-k^q\sin_{p,q}^q{\theta})
^{1/r}}\,d\theta
%&=\frac{q}{r^*k} \int_0^{\pi_{p,q}/2}
%\dfrac{1-k^q\sin_{p,q}^q{\theta}-1}{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}
%\,d\theta\\
=\frac{q}{r^*k} (E_{p,q,r}-K_{p,q,r}).
\end{align*}

Next, for $K_{p,q,r}(k)$ 
\begin{align*}
\frac{dK_{p,q,r}}{dk}
=\frac{q}{r}\int_0^{\pi_{p,q}/2}
\dfrac{k^{q-1}\sin_{p,q}^q{\theta}}
{(1-k^q\sin_{p,q}^q{\theta})^{1+1/r}}\,d\theta.
\end{align*}
Here we see that 
\begin{align*}
& \frac{d}{d\theta} 
\left(\frac{-\cos_{p,q}^{p/r}{\theta}}
{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}\right)
%&=\frac{(q/r)\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{1-p/r^*}{\theta}
%(1-k^q\sin_{p,q}^q{\theta})
%-(q/r)k^q\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{p/r+1}{\theta}}
%{(1-k^q\sin_{p,q}^q{\theta})^{1+1/r}}\\
=\frac{q(1-k^q)\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{1-p/r^*}{\theta}}
{r(1-k^q\sin_{p,q}^q{\theta})^{1+1/r}},\\
& \lim_{\theta \to \pi_{p,q}/2} \cos_{p,q}^{p-1}{\theta}
=\lim_{\theta \to \pi_{p,q}/2} (1-\sin_{p,q}^q{\theta})^{1/p^*}=0;
\end{align*}
so that we use integration by parts as
\begin{align*}
\frac{dK_{p,q,r}}{dk}
&=\frac{k^{q-1}}{1-k^q}\int_0^{\pi_{p,q}/2}
\frac{d}{d\theta} 
\left(\frac{-\cos_{p,q}^{p/r}{\theta}}
{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}\right)
\sin_{p,q}{\theta}\cos_{p,q}^{p/r^*-1}{\theta}\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \left[\frac{-\sin_{p,q}{\theta}\cos_{p,q}^{p-1}{\theta}}
{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}
\right]_0^{\pi_{p,q}/2}\\
& \qquad
+\frac{k^{q-1}}{1-k^q} \int_0^{\pi_{p,q}/2} 
\frac{\cos_{p,q}^{p/r}{\theta}}{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}
\left(\cos_{p,q}^{p/r^*}{\theta}-
%q\left(\frac{1}{r}-\frac{1}{p}\right)
\frac{(q/r^*-q/p)\sin_{p,q}^q{\theta}}{\cos_{p,q}^{p/r}{\theta}}\right)\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \int_0^{\pi_{p,q}/2} 
\frac{\cos_{p,q}^p{\theta}-(q/r^*-q/p)\sin_{p,q}^q{\theta}}
{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \int_0^{\pi_{p,q}/2}
\frac{(1+q/r^*-q/p)(1-k^q\sin_{p,q}^q{\theta})-(1+q/r^*-q/p-k^q)}{k^q(1-k^q\sin_{p,q}^q{\theta})^{1/r}}\,d\theta\\
&=\frac{(1+q/r^*-q/p)E_{p,q,r}-(1+q/r^*-q/p-k^q)K_{p,q,r}}{k(1-k^q)}.
\end{align*}

The case $p=\infty$ is proved similarly. Indeed,
\begin{align*}
\frac{dE_{\infty,q,r}}{dk}
%&=\int_0^{\pi_{p,q}/2}
%\frac{d}{dk}(1-k^q\sin_{p,q}^q{\theta})^{1/r^*}\,d\theta\\
&=\frac{q}{r^*}\int_0^1
\dfrac{-k^{q-1}\theta^q}{(1-k^q\theta^q)
^{1/r}}\,d\theta
%&=\frac{q}{r^*k} \int_0^{\pi_{p,q}/2}
%\dfrac{1-k^q\sin_{p,q}^q{\theta}-1}{(1-k^q\sin_{p,q}^q{\theta})^{1/r}}
%\,d\theta\\
=\frac{q}{r^*k} (E_{\infty,q,r}-K_{\infty,q,r})
\end{align*}
and
\begin{align*}
\frac{dK_{\infty,q,r}}{dk}
&=\frac{q}{r}\int_0^1
\dfrac{k^{q-1}\theta^q}
{(1-k^q\theta^q)^{1+1/r}}\,d\theta\\
%\end{align*}
%Here we see that 
%\begin{align*}
%& \frac{d}{d\theta} 
%\left(\frac{-\cos_{p,q}^{p/r^*}{\theta}}
%{(1-k^q\sin_{p,q}^q{\theta})^{1/r^*}}\right)
%&=\frac{(q/r)\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{1-p/r^*}{\theta}
%(1-k^q\sin_{p,q}^q{\theta})
%-(q/r)k^q\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{p/r+1}{\theta}}
%{(1-k^q\sin_{p,q}^q{\theta})^{1+1/r}}\\
%=\frac{q(1-k^q)\sin_{p,q}^{q-1}{\theta}\cos_{p,q}^{1-p/r}{\theta}}
%{r^*(1-k^q\sin_{p,q}^q{\theta})^{1+1/r^*}},\\
%& \lim_{\theta \to \pi_{p,q}/2} \cos_{p,q}^{p-1}{\theta}
%=\lim_{\theta \to \pi_{p,q}/2} (1-\sin_{p,q}^q{\theta})^{1/p^*}=0;
%\end{align*}
%so that we use integration by parts as
%\begin{align*}
%\frac{dK_{p,q,r}}{dk}
&=\frac{k^{q-1}}{1-k^q}\int_0^1
\frac{d}{d\theta} 
\left(-\left(\frac{1-\theta^q}{1-k^q\theta^q}\right)^{1/r}\right)
\theta(1-\theta^q)^{1/r^*}\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \left[\frac{-\theta(1-\theta^q)}
{(1-k^q\theta^q)^{1/r}}\right]_0^1\\
& \qquad
+\frac{k^{q-1}}{1-k^q} \int_0^1 
\left(\frac{1-\theta^q}{1-k^q\theta^q}\right)^{1/r}
\left((1-\theta^q)^{1/r^*}-\frac{(q/r)\theta^q}{(1-\theta^q)^{1/r}}\right)\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \int_0^1 
\frac{1-\theta^q-(q/r)\theta^q}
{(1-k^q\theta^q)^{1/r}}\,d\theta\\
&=\frac{k^{q-1}}{1-k^q} \int_0^1
\frac{(1+q/r^*)(1-k^q\theta^q)-(1+q/r^*-k^q)}{k^q(1-k^q\theta^q)^{1/r}}\,d\theta\\
&=\frac{(1+q/r^*)E_{\infty,q,r}-(1+q/r^*-k^q)K_{\infty,q,r}}{k(1-k^q)}.
\end{align*}
This completes the proof.
\end{proof}

%%%%%

Proposition \ref{prop:p-differential} now yields 
Theorem \ref{thm:pqrL}.

\begin{proof}[Proof of Theorem \ref{thm:pqrL}]
Let $k':=(1-k^q)^{1/r},\ E'_{p,r,q^*}(k):=E_{p,r,q^*}(k')$ and 
$K'_{p,r,q^*}(k):=K_{p,r,q^*}(k')$.
As $dk'/dk=-(q/r)k^{q-1}/(k')^{r-1}$,
Proposition \ref{prop:p-differential} gives
\begin{align*}
\frac{dE_{p,q,r^*}}{dk}&=\frac{q(E_{p,q,r^*}-K_{p,q,r^*})}{rk},\\
\frac{dK_{p,q,r^*}}{dk}&=\dfrac{aE_{p,q,r^*}-(a-k^q)K_{p,q,r^*}}{k(k')^r},\\
\frac{dE'_{p,r,q^*}}{dk}&=\frac{k^{q-1}(-E'_{p,r,q^*}+K'_{p,r,q^*})}{(k')^r},\\
\frac{dK'_{p,r,q^*}}{dk}&=\frac{q(-bE'_{p,r,q^*}+(b-(k')^r)K'_{p,r,q^*})}{rk(k')^r},
\end{align*}
where $a:=1+q/r-q/p$ and $b:=1+r/q-r/p$.

We denote the left-hand side of \eqref{eq:pqrL} by $L(k)$.
A direct computation shows that
\begin{align*}
\frac{d}{dk} & L(k)\\
&  = 
\frac{q(E_{p,q,r^*}-K_{p,q,r^*})}{rk} \cdot K'_{p,r,q^*}
+E_{p,q,r^*} \cdot \frac{q(-bE'_{p,r,q^*}+(b-(k')^r)K'_{p,r,q^*})}{rk(k')^r}\\
& \quad 
+\dfrac{aE_{p,q,r^*}-(a-k^q)K_{p,q,r^*}}{k(k')^r} \cdot E'_{p,r,q^*}
+K_{p,q,r^*} \cdot \frac{k^{q-1}(-E'_{p,r,q^*}+K'_{p,r,q^*})}{(k')^r} \\
& \qquad 
-\dfrac{aE_{p,q,r^*}-(a-k^q)K_{p,q,r^*}}{k(k')^r} \cdot K'_{p,r,q^*}
-K_{p,q,r^*} \cdot \frac{q(-bE'_{p,r,q^*}+(b-(k')^r)K'_{p,r,q^*})}{rk(k')^r}\\
& =
\left(\frac{q}{rk}+\frac{q(b-(k')^r)}{rk(k')^r}-\frac{a}{k(k')^r}\right)E_{p,q,r^*}K'_{p,r,q^*}\\
& \quad
+\left(-\frac{q}{rk}+\frac{k^{q-1}}{(k')^r}+\frac{a-k^q}{k(k')^r}-\frac{q(b-(k')^r)}{rk(k')^r}\right)K_{p,q,r^*}K'_{p,r,q^*}\\
& \qquad
+\left(-\frac{qb}{rk(k')^r}+\frac{a}{k(k')^r}\right)E_{p,q,r^*}E'_{p,r,q^*}\\
& \qquad \quad
+\left(-\frac{a-k^q}{k(k')^r}-\frac{k^{q-1}}{(k')^r}+\frac{qb}{rk(k')^r}\right)K_{p,q,r^*}E'_{p,r,q^*}\\
& =
\frac{qb-ra}{rk(k')^r}(E_{p,q,r^*}K'_{p,r,q^*}-K_{p,q,r^*}K'_{p,r,q^*}-E_{p,q,r^*}E'_{p,r,q^*}+K_{p,q,r^*}E'_{p,r,q^*}).
\end{align*}
Since $qb-ra=0$, we see that $dL/dk=0$.
Thus $L(k)$ is a constant $C$.

We will evaluate $C$ as follows. 
Since
\begin{align*}
|(K_{p,q,r^*} & -E_{p,q,r^*})K'_{p,r,q^*}| \\
&=\int_0^{\pi_{p,q}/2} 
\left(\frac{1}{(1-k^q\sin_{p,q}^q{\theta})^{1/r^*}}
-(1-k^q\sin_{p,q}^q{\theta})^{1/r}\right)\,d\theta\\
& \qquad \times \int_0^{\pi_{p,r}/2} 
\frac{d\theta}{(1-(k')^r\sin_{p,r}^r{\theta})^{1/q^*}}\\
&=\int_0^{\pi_{p,q}/2} 
\frac{k^q\sin_{p,q}^q{\theta}}{(1-k^q\sin_{p,q}^q{\theta})^{1/r^*}}\,d\theta \cdot
\int_0^{\pi_{p,r}/2} 
\frac{d\theta}{(\cos_{p,r}^p{\theta}+k^q\sin_{p,r}^r{\theta})^{1/q^*}}\\
& \leq k^q K_{p,q,r^*}(k) \cdot \frac{1}{k^{q-1}}\frac{\pi_{p,r}}{2}\\
& =\frac{\pi_{p,r}}{2}kK_{p,q,r^*}(k),
\end{align*}
we obtain $\lim_{k \to +0}(K_{p,q,r^*}-E_{p,q,r^*})K'_{p,r,q^*}=0$.
Therefore, from Proposition \ref{prop:behaviour} 
$$C=\lim_{k \to +0}K_{p,q,r^*}E'_{p,r,q^*}
=K_{p,q,r^*}(0)E_{p,r,q^*}(1)
=\frac{\pi_{p,q}\pi_{s,r}}{4},$$
%4r}B\left(\frac{1}{r},\frac{1}{p^*}+\frac{1}{q}\right).$$
where $1/s=1/p-1/q$. Thus, we conclude the assertion.
\end{proof}

Finally, we will give a remark for Theorem \ref{thm:pqrL}. 
From the series expansion and the termwise integration, 
it is possible to express the generalized complete elliptic integrals
by Gaussian hypergeometric functions
\begin{align*}
K_{p,q,r}(k)&=\frac{\pi_{p,q}}{2}F\left(\frac{1}{q},\frac{1}{r};\frac{1}{p^*}+\frac{1}{q};k^q\right),\\
E_{p,q,r}(k)&=\frac{\pi_{p,q}}{2}F\left(\frac{1}{q},-\frac{1}{r^*};\frac{1}{p^*}+\frac{1}{q};k^q\right).
\end{align*}
By these expressions and letting $1/p=1/2-b,\ 1/q=1/2+a,\ 1/r=1/2-c$ and $k^q=x$
in \eqref{eq:pqrL}, we obtain \textit{Elliott's identity}
(see Elliott \cite{El}; see also \cite{AQVV}, \cite[Theorem 3.2.8]{AAR} and 
\cite[(13) p.\,85]{EMOT}):
\begin{multline}
\label{eq:elliott}
F\left({{1/2+a,-1/2-c}\atop{a+b+1}} ;x\right)
F\left({{1/2-a,1/2+c}\atop{b+c+1}} ;1-x\right)\\
+F\left({{1/2+a,1/2-c}\atop{a+b+1}};x\right)
F\left({{-1/2-a,1/2+c}\atop{b+c+1}};1-x\right)\\
-F\left({{1/2+a,1/2-c}\atop{a+b+1}};x\right)
F\left({{1/2-a,1/2+c}\atop{b+c+1}};1-x\right)\\
=\frac{\Gamma(a+b+1)\Gamma(b+c+1)}{\Gamma(a+b+c+3/2)\Gamma(b+1/2)}
\end{multline}
for $|a|,\,|c|<1/2$ and $b \in (-1/2,\infty)$,
where $\Gamma$ denotes the gamma function.
Also, letting $1/p=2-c-a$ and $1/q=1-a$ in \eqref{eq:pqL} of 
Corollary \ref{cor:q=r}, we have the identity of \cite[Corollary 3.13 (5)]{AQVV}
for $a \in (0,1)$ and $c \in (1-a,\infty)$.
A series of Vuorinen's works on Elliott's identity with his coauthors 
starting from \cite{AQVV}
deals with the concavity/convexity properties of certain related functions
to the left-hand side of \eqref{eq:elliott}.

%For $a=b=c=0$, \eqref{eq:elliott} coincides with Legendre's relation \eqref{eq:L}.

%\begin{multline}
%\label{eq:elliott}
%F\left({{\frac12+a,-\frac12-c}\atop{a+b+1}} ;x\right)
%F\left({{\frac12-a,c+\frac12}\atop{b+c+1}} ;1-x\right)\\
%+F\left({{a+\frac12,\frac12-c}\atop{a+b+1}};x\right)
%F\left({{-(a+\frac12),c+\frac12}\atop{b+c+1}};1-x\right)\\
%-F\left({{a+\frac12,\frac12-c}\atop{a+b+1}};x\right)
%F\left({{\frac12-a,c+\frac12}\atop{b+c+1}};1-x\right)\\
%=\frac{\Gamma(a+b+1)\Gamma(b+c+1)}{\Gamma(a+b+c+\frac32)\Gamma(b+\frac12)},
%\end{multline}
%Of course, from Proposition \ref{prop:hypergeometricexpression} and Elliott's
%identity with $a=-b=c=1/p-1/2$ and $x=k^p$,
%we are also able to obtain Theorem \ref{thm:p-legendre}.

%%%%

% \newpage

% Abbreviated version of the title for the running head:

% ``Generalized elliptic functions and its applications to
% the $p$-Laplacian''
% \bigskip\\

% Name and mailing address of the author to whom proofs should be sent:

% Shingo Takeuchi

% shingo@cc.kogakuin.ac.jp

\end{document}